OpenStudy (anonymous):

The graph of y=4^x-1 shifts the graph of y=4^x _______ . A. left 1 unit B. right 1 unit C. up 1 unit D. down 1 unit

OpenStudy (skyz):

I think D. o-o I can't be sure cuz I've never seen a question asked liek this about a graph

OpenStudy (mathstudent55):

Is it \(y = 4^x - 1\) or \(y = 4^{x-1} \) compared to \(y = 4^x\) ? Each of those is a proper question and will have a different answer.

OpenStudy (mathstudent55):

If it is \(y = 4^{x - 1} \) compared to \(y = 4^x\), then when you replace x with x - k, the graph shifts k units to the right if k is positive and k units to the left if k is negative.

OpenStudy (mathstudent55):

If it is \(y = 4^x - 1 \) compared to \(y = 4^x\), then it's the same as \(y + 1 = 4^x\). When you replace y with y - k, the graph shifts k units up if k is positive and k units down if k is negative.